# Biparental Reproduction may Enhance Species Sustainability by Conserving Shared Parental Traits more Faithfully than Monoparental Reproduction

### v2

@@ Abstract
Recognized effects of sexual reproduction and other forms of biparental reproduction in species sustainment and evolution include the increasing of diversity, accelerating adaptation, constraining the accumulation of deleterious mutations, and the homogenization of species genotype. Nevertheless, many questions remain open with regard to the evolution of biparental reproduction. In this paper we contribute an initial exploration of finer details of the homogenization effect that we believe deserve focused analysis and discussion. Specifically, rudimentary mathematical analyses suggest the perspective that biparental reproduction enhances the retention in the offspring of shared, i.e., common, properties of the parent generation, as compared with monoparental, clonal reproduction. We argue that this is an intrinsic effect of merging the encodings of parents’ traits, regardless of physical, chemical, biological and social aspects of the reproduction process and of the trait at hand. As the survival of offspring often depends on their ability to step in and actively fill the voids left by failing, dying or absent members of the species in interaction networks that sustain the species, cross-generation retention of common species properties helps sustain the species. This side-effect of sexual reproduction may have also contributed to the very evolution and pervasiveness of sexual reproduction in nature.
@@ 1 Introduction
Natural reproduction is of course about transferring parental features to offspring. It is also well accepted that biparental reproduction yields resemblance of offspring in certain aspects to one parent or the other, or to both, or to having a mixture of traits from both parents. In this paper we explore a finer effect of sexual reproduction which is not addressed explicitly in discussions of evolution of sexual reproduction like those cited below, namely, that biparental reproduction conserves traits that are shared by both parents, and that may be common to the entire parent generation within the species.
@@
While there are multiple theories for the sustainment and evolutionary effects of biparental reproduction, both sexual (with two or more mating types) and unisexual (with one mating type, as is common, for example in fungi [Heitman, 2015]), over monoparental, clonal, asexual reproduction, the underlying question is still considered open [MacPherson et al., 2023, MacPherson et al., 2021, Livnat and Papadimitriou, 2016, Billiard et al., 2012]. and others.
@@
The most commonly cited consequences of biparental reproduction include: (i) the introduction of diversity, which then leads to increased adaptivity to changing conditions, and enhanced competitiveness and the ability to co-evolve with pathogens, parasites, or symbionts; for example the Red Queen hypothesis [Van Valen, 1973]; (ii) the constraining the irreversible penetration of deleterious mutations, often termed Muller’s ratchet [Kondrashov, 1988, Muller, 1932]; (iii) preservation of common properties by facilitating the raising of reproduction barriers between divergent subgroups within a species [Barton et al., 2007, CH.22], DNA repair [Bernstein et al., 1981], [Barton et al., 2007, Ch.12 p.335] that is part of meiosis, sexual selection which promotes reproduction of individuals with certain preferred properties, and interference with specialization and speciation of variants that find and prefer new niches [Barton et al., 2007, Ch.16]. Indeed, the genome homogenization effect of biparental reproduction is often taken for granted when a population is panmictic, i.e., every organism of the species can mate with every other of the same species, and confined within a geographical region [Fitzpatrick et al., 2008, Abstract]; it is also often implicit from discussions of gene flow in structured populations [Barton et al., 2007, Ch.16 p.441.], or in pondering the puzzle of sympatric speciation, i.e., why do panmictic populations in a given location sometime give rise to diversity and speciation [Fitzpatrick et al., 2008, Mayr, 1982].
@@
Here we show that in a population whose clonal reproduction is imperfect and may induce some change, as is the case with many kinds of DNA mutations [Barton et al., 2007, Ch.12], horizontal gene transfer [Keeling and Palmer, 2008] incorporation of viral genetic material [Koonin, 2010], etc., biparental reproduction that employs merging of encodings of inherited properties, has an underlying mathematical conservation effect that is independent of biological specifics, and is more general than the homogenization mechanisms listed above. Specifically, biparental reproduction conserves the properties that are common within the parents’ generation better than monoparental reproduction. This, of course, then extends to conservation of ancestral properties across multiple generations.
@@
We support our claim that the effects of biparental reproduction are intrinsic, by initially showing that they exist in an abstracted setting: we deal with only a single property, and the mechanism does not attempt to model in detail meiosis, genetic recombination, epistasis, syngamy, the role of mating types, the role of the environment, and other complex aspects of reproduction in nature. While all these factors are of course important and should be part of realistic models of reproduction in nature, the particular aspects of biparental reproduction discussed here deserve a special focus beyond the published theories, mechanisms and effects.
@@
    Consider at first the prominent effect of sexual reproduction, which is the enhancing species diversity. This capability relies on the fact that individuals differ from each other in the first place. By contrast, the effect discussed here focuses on the traits that do not change
@@
    The way sexual reproduction and associated gene flow constrains specialization and fitting into new niches by diluting new beneficial traits may be a special case of the mechanism described here. Biparental reproduction slows down property drift regardless of how useful the new properties may be in new niches. We emphasize that interference with speciation comes from retention of shared traits at the expense of diverse ones. Hypothetically the dilution of over-specialized properties could occur by further randomizing then, and preventing them from becoming common and shared, without slowing down the drifting away of the properties that are shared and common.
@@
    Most importantly, homogenization of the next generation as induced by sexual reproduction, and discussed extensively in the literature, does not necessarily imply that a shared property of a homogeneous new generation is also closer to the shared property of the parent generation. For example, if the parent generation is comprised of clusters of different trait values, the child generation may have genotype and phenotype properties that are not in any of these clusters.
In the next section, we briefly discuss some of the motivation for the particular principles applied in the ensuing model, and in Section 3 we use a synthetic model with encodings consisting of sets of real numbers to prove mathematically the inevitability of conservation of parental properties as a general effect of biparental reproduction derived from basic probability and the central limit theorem. In future versions we plan to add numerical examples and simulation runs. Version 1 of the arXiv preprint contains such numerical examples, and analysis and visualization of simulation runs of a slightly different model.
In summary, given that survival of organisms, species and entire ecosystems depends on organisms fitting in into a rich interaction network, and given that all living individuals are destined to eventually die, our interim results show intrinsic properties of biparental reproduction likely help sustain each species by enabling new generations to carry out the interaction roles performed by their parents and by several preceding generations. Further studies are needed to generalize and expand the model, and to confirm that the conservation of shared traits by biparental reproduction is indeed a distinct natural phenomenon.
@@ 2 Modeling Rationale
Before delving into the details of the mathematical model, in this section we offer some background and context for various choices applied in the model and its analysis. The reader may choose to skip this section, and return to it as needed.
2.1 Codes, Interaction codes and Autoencoding
@@ 2.2 Information aggregation by set operations
The present setting is amenable to extensions using more elaborate methods for deriving child traits from parents’
Note that pure averaging of a single property represented by a random variable would be analogous to combining two one-dimensional random walks, with the same step size and equal probabilities for stepping in either direction. Such random walks can be readily shown to have a mean position which is the average of the two origins, and, more importantly, its variance is half the (equal) variance of the two original random walks. This is aligned with the known fact that biparental reproduction reduces the variance within the offspring generation.
@@ 2.3 Minimalism, abstraction and accessible language
The model is not intended to be a full model of what happens in nature. In fact, we try to show that in situations of noisy inheritance, it is the very merging of parent encodings that enables the cross-generation conservation of ancestral traits. Examples of elements of the abstraction that we employed include: (i) The property of interest and its encoding both involve real numbers. Extending the arguments to pure symbolic entities, or concrete objects like molecules, is left for a future research; (ii) The inevitable variation in natural reproduction is modeled by replacing direct copying with sampling from a normal distribution; (iii) the biparental reproduction is unisexual, i.e., there are no distinct mating types.
@@
3 Problem formalization and claims
In this section we compare biparental and monoparental reproduction in the context of a specific model.
3.1 The underlying model
@@
In the present discussion it does not matter whether the parent generation persists or disappears as long as the entities do not reproduce. For simplicity we assume that the entire cohort of each parent generation disappears immediately and synchronously upon the appearance of the cohort of the child generation. One reason for this choice is to avoid dealing with inter-generation mating, in a population in which diversity among co-existing distant generations is unbounded.
@@
The “genotype” encoding of $P$ in each individual $e_{i}$ is a set $g_{i}$ of an even number $k$ of real values which are also referred to as the points in $g_{i}$. We allow the rare case of any two points in a set being equal to each other, without resorting to other mathematical terms like $g_{i}$ being a multiset or a tuple. The integer $k$ is a fixed parameter in the model, and it is large, say, $k\geq 100$.
@@ 3.2 Reproduction methods
We model and compare the effects of two different reproduction processes, termed Monoparental Reproduction (MR, hereafter) and Biparental Reproduction (BR, hereafter).
@@
      Create two interim (“gamete”) encodings $g^{\prime}_{u}$ and $g^{\prime}_{v}$ by randomly selecting $k/2$ points, under normal distribution with mean $p_{u}$ and variance $h^{2}$, and $k/2$ points under normal distribution with mean $p_{v}$ and variance $h^{2}$, respectively; let $g^{\prime}$ be the union of these two sets, $g^{\prime}=g^{\prime}_{u}\cup g^{\prime}_{v}$. We term this process merging of the parents’ encodings. Note that this merging involves a sampling step, and does not rely on copying of elements of $g_{u}$ or $g_{v}$ into $g^{\prime}$.
@@
3.3 Analytical claims
In comparing MR and BR in a particular transition from a parent cohort $c$ to a child cohort $c^{\prime}$, we are initially interested in a basic metric $d$ defined as the distance between the phenotype properties of a random individual in $c^{\prime}$, denoted $p^{\prime}_{i}$, and the common property $p_{c}$ of the parents’ cohort
$d_{i}=|p^{\prime}_{i}-p_{c}|$
@@ 3.3 Analytical claims
The genotype encodings $g^{\prime}_{i_{\text{MR}}}$ of all possible direct MR children of $e_{i}$, can be considered as samples of size $k$ from a normal distribution with mean $p_{i}$ and variance $h^{2}$.
According to the central limit theorem the means of these samples, namely $p^{\prime}_{i_{\text{MR}}}$, are distributed normally with mean $p_{i}$ and standard deviation $\frac{h}{\sqrt{k}}$ (and variance $\frac{h^{2}}{k}$).
We are facing now a compound, nested distribution, that involves (i) the random selection of an $e_{i}$, which is close enough for our purposes to drawing a random $p_{i}$ from the original normal distribution of $c$ with mean $p_{c}$ and variance $\sigma^{2}_{c}$, and (ii) based on the selected $p_{i}$ and the noise parameter, randomly selecting the sample $g^{\prime}$ of size $k$ and computing its mean $p^{\prime}_{i_{\text{MR}}}$.
@@
The genotype encoding $g^{\prime}_{j_{\text{BR}}}$ was created as a union of two random samples $g^{\prime}_{u}$ and $g^{\prime}_{v}$, each of size $k/2$, around the means $p_{u}$ and $p_{v}$, respectively. Let $p^{\prime}_{u}$ and $p^{\prime}_{v}$ be the means of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, respectively.
@@
The sets $g^{\prime}_{u}$ and $g^{\prime}_{v}$ were drawn from distributions with means $p_{u}$ and $p_{v}$ respectively, both with variance $h^{2}$. According to the Central Limit Theorem again, the distributions of $p^{\prime}_{u}$ and $p^{\prime}_{v}$ are normal with mean $p_{u}$ and $p_{v}$ respectively, and standard deviation $\frac{h}{\sqrt{\frac{k}{2}}}$ (and variance $\frac{h^{2}}{\frac{k}{2}}$).
@@
Clearly the mean of the random variable $p^{\prime}_{j_{\text{BR}}}$ is $p_{c}$ since it is half of the sum $E[p^{\prime}_{u}]$ and $E[p^{\prime}_{v}]$, and its variance of is a quarter of the variance of their sum, as it is is the square of the scaling of the variable and the mean:
@@
Thus, $d_{\text{BR}}<d_{\text{MR}}$. $\square$

### v3

# Bi-parental Reproduction may Enhance Species Sustainability by Conserving Shared Parental Traits more Faithfully than Mono-parental Reproduction

@@ Abstract
There are many recognized effects of sexual reproduction and other forms of bi-parental reproduction on the sustainment and evolution of species. They include increased diversity, accelerated adaptation, constrained accumulation of deleterious mutations, and the homogenization of species’ genotype. Nevertheless, many questions remain open regarding the evolution of bi-parental reproduction. Here we explore some of the finer details of the homogenization effect, which we believe deserve deeper analysis and discussion. We present a first-cut mathematical analysis, which suggests that bi-parental reproduction enhances the retention of shared, common properties of the parent generation in their offspring, as compared with mono-parental, clonal reproduction. For illustration, we accompany the mathematical claim and proof with several simulation examples depicting and comparing bi- and mono-parental reproduction of real numbers, images of printed text, and images of flowers. We argue that the retention of shared traits of preceding generations is an intrinsic effect of merging the encodings of parents’ traits, regardless of physical, chemical, biological and social aspects of the reproduction process and of the specific trait at hand. As survival of offspring often depends on their ability to step in and fill voids left by failing, dying or absent members of the species in vital interaction networks, retention of common properties across generations helps sustain the species. This side-effect of sexual reproduction may have also contributed to the very evolution and pervasiveness of sexual reproduction in nature.
@@ 1 Introduction
Natural reproduction is of course about transferring parental features to offspring. It is also well accepted that bi-parental reproduction yields resemblance of offspring in certain aspects to one parent or the other, or to both, or to having a mixture of traits from both parents. In this paper we explore a finer effect of sexual reproduction which is not addressed explicitly in discussions of evolution of sexual reproduction like those cited below, namely, that bi-parental reproduction conserves traits that are shared by both parents, and that may be common to the entire parent generation within the species.
@@
While there are multiple theories for the sustainment and evolutionary effects of bi-parental reproduction, both sexual (with two or more mating types) and unisexual (with one mating type, as is common, for example in fungi [Heitman, 2015]), over mono-parental, clonal, asexual reproduction, the underlying question is still considered open [MacPherson et al., 2023, MacPherson et al., 2021, Livnat and Papadimitriou, 2016, Billiard et al., 2012]. and others.
@@
The most commonly cited consequences of bi-parental reproduction include: (i) the introduction of diversity, which then leads to increased adaptivity to changing conditions, and enhanced competitiveness and the ability to co-evolve with pathogens, parasites, or symbionts; for example the Red Queen hypothesis [Van Valen, 1973]; (ii) the constraining the irreversible penetration of deleterious mutations, often termed Muller’s ratchet [Kondrashov, 1988, Muller, 1932]; (iii) preservation of common properties by facilitating the raising of reproduction barriers between divergent subgroups within a species [Barton et al., 2007, CH.22], DNA repair [Bernstein et al., 1981], [Barton et al., 2007, Ch.12 p.335] that is part of meiosis, sexual selection which promotes reproduction of individuals with certain preferred properties, and interference with specialization and speciation of variants that find and prefer new niches [Barton et al., 2007, Ch.16]. Indeed, the genome homogenization effect of bi-parental reproduction is often taken for granted when a population is panmictic, i.e., every organism of the species can mate with every other of the same species, and confined within a geographical region [Fitzpatrick et al., 2008, Abstract]; it is also often implicit from discussions of gene flow in structured populations [Barton et al., 2007, Ch.16 p.441.], or in pondering the puzzle of sympatric speciation, i.e., why do panmictic populations in a given location sometime give rise to diversity and speciation [Fitzpatrick et al., 2008, Mayr, 1982].
@@
Here we show that in a population whose clonal reproduction is imperfect and may induce some change, as is the case with many kinds of DNA mutations [Barton et al., 2007, Ch.12], horizontal gene transfer [Keeling and Palmer, 2008] incorporation of viral genetic material [Koonin, 2010], etc., bi-parental reproduction that employs merging of encodings of inherited properties, has an underlying mathematical conservation effect that is independent of biological specifics, and is more general than the homogenization mechanisms listed above. Specifically, bi-parental reproduction conserves the properties that are common within the parents’ generation better than mono-parental reproduction. This, of course, then extends to conservation of ancestral properties across multiple generations. This view can be aligned with the observation in [Cohen, 2016] that one of the effects of sexual reproduction is “to maintain a frequency distribution of genes within the species that fits the life style and ecosystem”.
@@
We support our claim that the effects of bi-parental reproduction are intrinsic, by initially showing that they exist in an abstracted setting: we deal with only a single property, and the mechanism does not attempt to model in detail meiosis, genetic recombination, epistasis, syngamy, the role of mating types, the role of the environment, and other complex aspects of reproduction in nature. While all these factors are of course important and should be part of realistic models of reproduction in nature, the particular aspects of bi-parental reproduction discussed here deserve a special focus beyond the published theories, mechanisms and effects.
@@
    Consider at first the prominent effect of sexual reproduction, which is the enhancing of species diversity. This capability relies on the fact that individuals differ from each other in the first place. By contrast, the effect discussed here focuses on the traits that do not change
@@
    The way sexual reproduction and associated gene flow constrains specialization and fitting into new niches by diluting new beneficial traits may be a special case of the mechanism described here. Bi-parental reproduction slows down property drift regardless of how useful the new properties may be in new niches. We emphasize that interference with speciation comes from retention of shared traits at the expense of diverse ones. Hypothetically, the dilution of over-specialized properties could occur by further randomizing then, and preventing them from becoming common and shared, without slowing down the drifting away of the properties that are shared and common.
@@
    Most importantly, homogenization of the next generation as induced by sexual reproduction, and discussed extensively in the literature, does not necessarily imply that a shared property of a homogeneous new generation is also closer to the shared property of the immediately preceding parent generation. For example, if the parent generation is comprised of clusters of different trait values, a homogeneous child generation may still have genotype and phenotype properties that are not in any of these clusters.
In the next section, we briefly discuss some of the motivation for the particular principles applied in the ensuing model. In Section 3 we use a synthetic model with encodings consisting of sets of real numbers to prove mathematically the inevitability of conservation of parental properties as a general effect of bi-parental reproduction derived from basic probability and the central limit theorem. In Section 5 we illustrate the effect with examples drawn from simulation runs that reproduce diverse objects including abstract sets of points, images of printed text, and images of flowers.
In summary, given that survival of organisms, species and entire ecosystems depends on organisms fitting in into a rich interaction network, and given that all living individuals are destined to eventually die, our interim results show that intrinsic properties of bi-parental reproduction likely help sustain each species by enabling new generations to carry out the interaction roles performed by their parents and by several preceding generations. Further studies are needed to generalize and expand the model, and to confirm that the conservation of shared traits by bi-parental reproduction is indeed a distinct natural phenomenon.
@@ 2 Modeling Rationale
Before delving into the details of the mathematical model and the illustrating simulations, in this section we offer some background and context for various choices applied in the model and its analysis. The reader may choose to skip this section, and return to it as needed.
2.1 Codes, interaction codes, and autoencoding
@@ 2.2 Information aggregation by set operations
The present setting is amenable to extensions using more elaborate methods for deriving child traits from parents’ traits.
Note that pure averaging of a single property represented by a random variable would be analogous to combining two one-dimensional random walks, with the same step size and equal probabilities for stepping in either direction. Such random walks can be readily shown to have a mean position which is the average of the two origins, and, more importantly, its variance is half the (equal) variance of the two original random walks. This is aligned with the known fact that bi-parental reproduction reduces the variance within the offspring generation.
@@ 2.3 Minimalism, abstraction and accessible language
The model is not intended to be a full model of what happens in nature. In fact, we try to show that in situations of noisy inheritance, it is the very merging of parent encodings that enables the cross-generation conservation of ancestral traits. Examples of elements of the abstraction that we employed include: (i) The property of interest and its encoding both involve real numbers. Extending the arguments to pure symbolic entities, or concrete objects like molecules, is left for a future research; (ii) The inevitable variation in natural reproduction is modeled by replacing direct copying with sampling from a normal distribution; (iii) the bi-parental reproduction is unisexual, i.e., there are no distinct mating types.
@@
3 Problem Formalization
In this section we define the bi-parental and mono-parental reproduction methods in the context of a specific model.
3.1 The formal model
@@
In the present discussion it does not matter whether the parent generation persists or disappears as long as the entities in the parent generation do not reproduce after the appearance of the child generation. For simplicity, we assume that the entire cohort of each parent generation disappears immediately and synchronously upon the appearance of the cohort of the child generation. One reason for this choice is to avoid dealing with inter-generation mating in a population in which diversity among co-existing distant generations is unbounded.
@@
The “genotype” encoding of $P$ in each individual $e_{i}$ is a set $g_{i}$ of $k$ of real values which are also referred to as the points in $g_{i}$. We allow the rare case of any two points in a set being equal to each other, without resorting to other mathematical terms like $g_{i}$ being a multiset or a tuple. The integer $k$ is a fixed parameter in the model, and it is large, say, $k\geq 100$; we also constrain $k$ to be an even number.
@@ 3.2 Reproduction methods
We model and compare the effects of two different reproduction processes, termed Mono-parental Reproduction (MR, hereafter) and Bi-parental Reproduction (BR, hereafter).
@@
      Create two interim (“gamete”) encodings $g^{\prime}_{u}$ and $g^{\prime}_{v}$ by randomly selecting $\frac{k}{2}$ points, under normal distribution with mean $p_{u}$ and variance $h^{2}$, and $\frac{k}{2}$ points under normal distribution with mean $p_{v}$ and variance $h^{2}$, respectively; let $g^{\prime}$ be the union of these two sets, $g^{\prime}=g^{\prime}_{u}\cup g^{\prime}_{v}$. We term this process merging of the parents’ encodings. Note that this merging involves a sampling step, and does not rely on copying of elements of $g_{u}$ or $g_{v}$ into $g^{\prime}$.
@@
3.3 Model realization example
Below we illustrate the above formal model with an example realization in an imaginary species of intelligent and capable animals. One of the traits that sustains this species is their ability to build every year a mud hill of “exactly” a certain height, as traps for some creature that they feed upon. The way they remember the right height is that each mature individual carries with them a bunch of sticks that they got from their parents. The individual builds each mud hill so that its height is close to the mean length of its set of sticks. Old mud hills from the previous year are not available for comparison, as they have been washed away by the rains. When young individuals mature, they leave their parents to look for new territory. Before they leave, the parents (one or two of them, as the case may be) equip the young with a new bunch of sticks, cut as best as they can to the mean length of the parents’ bunch. In the case of two parents’ each one gives the child half the number of sticks based on that parent’s set. The parents cannot give their own sticks to the offspring, as they have to keep their own sticks for several more years. The individuals are solitary most of the time, so each one needs their own set of sticks. Individuals of the species use sets of sticks for this purpose, rather than just one, as the sticks may break or get lost. In addition, the parent’s skill and the available material limit the precision of each cut stick; having several sticks, some too short and some too long, is a way for conveying the height of the mud hill that is desired by the individual preparing the new set.
3.4 A numerical example
For illustration, below we go through the core elements of MR and BR using a small-scale example.
  1. 1.
    Parameter Setting: Assume that in the parents’ cohort $c$ , $P$ is normally distributed with mean $p_{c}=2.0$, with standard deviation $\sigma_{c}=0.2$ The number of points in a genotype is $k=4$; the points in a genotype set are shown here in a sorted order for easier reading. The standard deviation of the normal distribution of the reproduction noise is $h=0.15$
  2. 2.
    Let individual $e_{1}$ have a genotype encoding of $g_{1}=\{1.70,1.80,1.90,2.10\}$ whose mean and phenotype is $p_{1}=1.875$;
  3. 3.
    MR of $e_{1}$: Compute the mean of the genotype, yielding again $p_{1}=1.875$. Draw a sample of $k=4$ points around this value as a mean with the above standard deviation $h=0.15$, yielding say, $g^{\prime}=\{1.61,1.71.1.91,2.01\}$, whose mean (and phenotype of child $e_{\text{1,MR}}^{\prime}$) is p’=1.81;
  4. 4.
    Let individual $e_{2}$ in the parent cohort $c$ have genotype encoding of $g_{2}~{}=\{1.95,2.15,2.25,2.35\}$. whose mean and phenotype is $p_{2}=2.175$.
  5. 5.
    BR of $e_{1}$ and $e_{2}$: We sample $\frac{k}{2}=2$ points around the mean $p_{1}$, and $2$ points around the mean $p_{2}$, both with the noise standard deviation $h$, yielding, say, $g^{\prime}_{1}=\{1.8,2.1\}$, and $g^{\prime}_{2}=\{2.02,2.18\}$, respectively. The union of these two samples yields $g^{\prime}=\{{\color[rgb]{0,0,1}1.8},{\color[rgb]{1,0,0}2.02},{\color[rgb]{0,0,1}2.1},{\color[rgb]{1,0,0}2.18}\}$, whose mean (and phenotype of the joint offspring $e_{\text{1,2,BR}}^{\prime}$) is $p^{\prime}=2.025$.
4 Comparing the Two Reproduction Methods
In comparing MR and BR in a particular transition from a parent cohort $c$ to a child cohort $c^{\prime}$, we are initially interested in a basic metric $d$ defined as the distance between the phenotype properties of a random individual in $c^{\prime}$, denoted $p^{\prime}_{i}$, and the common property $p_{c}$ of the parents’ cohort:
$d_{i}=|p^{\prime}_{i}-p_{c}|~{}.$
@@ 4 Comparing the Two Reproduction Methods
The genotype encodings $g^{\prime}_{i_{\text{MR}}}$ of all possible direct MR children of $e_{i}$, can be considered as samples of size $k$ from a normal distribution with mean $p_{i}$ and variance $h^{2}$, where $h$ is the noise parameter defined above.
According to the central limit theorem [Blitzstein and Hwang, 2015, p.435], the means of these samples, namely $p^{\prime}_{i_{\text{MR}}}$, are distributed normally with mean $p_{i}$ and standard deviation $\frac{h}{\sqrt{k}}$ (and variance $\frac{h^{2}}{k}$).
We are facing now a compound, nested distribution, that involves (i) the random selection of an $e_{i}$, which is close enough for our purposes to drawing a random value $p_{i}$ from the original normal distribution of $c$ with mean $p_{c}$ and variance $\sigma^{2}_{c}$, and (ii) based on the selected $p_{i}$ and the noise parameter, randomly selecting the sample $g^{\prime}$ of size $k$ and computing its mean $p^{\prime}_{i_{\text{MR}}}$.
@@
The genotype encoding $g^{\prime}_{j_{\text{BR}}}$ was created as a union of two random samples $g^{\prime}_{u}$ and $g^{\prime}_{v}$, each of size $\frac{k}{2}$, around the means $p_{u}$ and $p_{v}$, respectively. Let $p^{\prime}_{u}$ and $p^{\prime}_{v}$ be the means of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, respectively.
@@
The sets $g^{\prime}_{u}$ and $g^{\prime}_{v}$ were drawn from distributions with means $p_{u}$ and $p_{v}$ respectively, both with variance $h^{2}$. According to the central limit theorem again, the distributions of $p^{\prime}_{u}$ and $p^{\prime}_{v}$ are normal with mean $p_{u}$ and $p_{v}$ respectively, and standard deviation $\frac{h}{\sqrt{\frac{k}{2}}}$ (and variance $\frac{h^{2}}{(\frac{k}{2})}$).
@@
Clearly the mean of the random variable $p^{\prime}_{j_{\text{BR}}}$ is $p_{c}$ since it is half of the sum $E[p^{\prime}_{u}]$ and $E[p^{\prime}_{v}]$ (each of which is equal to $p_{c}$), and its variance is a quarter of the variance of their sum, as it is the square of the scaling by half of the variable and the mean [Blitzstein and Hwang, 2015, p.159]:
@@
Since the first term in the expression for $d_{\text{BR}}$ is $\frac{\sigma^{2}_{c}}{2}$ and for $d_{\text{MR}}$ it is $\sigma^{2}_{c}$, we get that
$d_{\text{BR}}<d_{\text{MR}}~{}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\square$